OpenStudy (aryana_maria2323):

Find dy, dx for the square root of y equals 3 plus the square root of x. a. y divided by x b. the square root of y divided by the square root of x c. the square root of x divided by the square root of y d. x divided by y

OpenStudy (mathmale):

Please type out your functions if at all possible. Is the following correct?\[y=3+\sqrt{x}\]

OpenStudy (mathmale):

1. What is the derivative of a constant, such as 3? 2. What is the exponential form of Sqrt(x)? 3. What is the derivative of \[y=x ^{1/2}\]

OpenStudy (aryana_maria2323):

dy/dx for \[\sqrt{y}= 3 + \sqrt{x}\]

OpenStudy (mathmale):

Convert both square roots to exponential form, please.

OpenStudy (aryana_maria2323):

The thing that is throwing me off is all the square roots.

OpenStudy (mathmale):

2. What is the exponential form of Sqrt(x)? 3. What is the derivative of y=x^(1/2) ?

OpenStudy (mathmale):